OpenStudy (anonymous):

Find the vertex, axis of symmetry y=4x^2-16x+3 Find the vertex, axis of symmetry y=4x^2-16x+3 @Mathematics

OpenStudy (anonymous):

For the vertex, x - value of vertex = -b / 2a = -(-16) / 2(4) = 2 y -value of vertex = 4(2)^2 - 16(2) + 3 = 16 - 32 + 3 = -13 vertex is at (2, -13)

OpenStudy (anonymous):

Axis of symmetry is x = 2

OpenStudy (anonymous):

what would be the x and y intercepts?

OpenStudy (anonymous):

x intercepts are the zeroes. x = 3.8 and x = 0.19 The y-intercept is when x = 0. y = 4(0)^2 - 16(0) + 3 = 0 - 0 +3 = 3

OpenStudy (anonymous):

You can round 0.19 to 0.2.